SPS PP) 0.5 0 0.1 0.0 0.4 P(P 0.3 0.1 0.2 0.8 Above table shows probabilities regarding customers soda and popcorn purchases at a movie theater. 13-What is the probability a randomly selected customer will purchase soda, L.e. P(S)? 14-What is the probability a randomly selected customer will purchase popcorn, i.e. P(P)? 15-What is the probability a randomly selected customer will not purchase soda? 16-What is the probability a randomly selected customer will not purchase popcorn? 17-What is the probability a randomly selected customer will purchase popcorn and soda together? 18-What is the probability a randomly selected customer will purchase popcorn or soda (or both)? 19-What is the probability a randomly selected customer purchases soda given that he purchases popcorn? 20-What is the probability a randomly selected customer purchases popcorn given that he purchases soda
Answer #1

Question 13

Pr(A randomly selected student will purchase soda) = P(s) = 0.5 + 0.3 = 0.8

Question 14

Pr(Will purcahse popcorn) = P(P) = 0.5 + 0.1 = 0.6

Question 15

Pr(Will not purchase soda) = 1- P(s) = P(sc) = 1 - 0.8 = 0.2

Question 16

Pr(Will not purchase popcorn) = 1 - P(P) = 1 - 0.6 = 0.4

Question 17

Pr(Will purchase popcorn and soda together) = 0.5 (as it is the common probability)

Question 18

Pr(Will purchase popcorn and soda or both) = 0.5 +  0.3 + 0.1 = 0.9

Question 19

Pr(Purchases soda given that he purchase popcorn) = 0.5/(0.5 + 0.1) = 0.8333

Question 20

Pr(Purchase popcorn given that her purchase soda) = 0.5/(0.5 + 0.3) = 0.5/0.8 = 0.625